### v2

# HeLo:::: Heterogeneous Multi-Modal Fusion with Label Correlation for Emotion Distribution Learning

Abstract.
Multi-modal emotion recognition has garnered increasing attention as it plays a significant role in human-computer interaction (HCI) in recent years. Since different discrete emotions may exist at the same time, compared with single-class emotion recognition, emotion distribution learning (EDL) that identifies a mixture of basic emotions has gradually emerged as a trend. However, existing EDL methods face challenges in mining the heterogeneity among multiple modalities. Besides, rich semantic correlations across arbitrary basic emotions are not fully exploited. In this paper, we propose a multi-modal emotion distribution learning framework, named HeLo, aimed at fully exploring the heterogeneity and complementary information in multi-modal emotional data and label correlation within mixed basic emotions. Specifically, we first adopt cross-attention to effectively fuse the physiological data. Then, an optimal transport (OT)-based heterogeneity mining module is devised to mine the interaction and heterogeneity between the physiological and behavioral representations. To facilitate label correlation learning, we introduce a learnable label embedding optimized by correlation matrix alignment. Finally, the learnable label embeddings and label correlation matrices are integrated with the multi-modal representations through a novel label correlation-driven cross-attention mechanism for accurate emotion distribution learning. Experimental results on two publicly available datasets demonstrate the superiority of our proposed method in emotion distribution learning.
Emotion Recognition, Emotion Distribution Learning, Label Correlation, Optimal Transport
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Human-centered computing HCI design and evaluation methodsccs: Computing methodologies Artificial intelligence
1. Introduction
Emotion recognition plays a pivotal role in enhancing human-computer interaction (HCI) (Shneiderman and Plaisant, 2010; Li et al., 2023; Can et al., 2023), offering substantial benefits across various fields such as healthcare, education, and customer service (Alarcao and Fonseca, 2017). Researchers have investigated various behavioral signals, including facial expressions (Verma et al., 2021) and voice signals (Yi and Mak, 2020), as well as physiological signals like electroencephalogram (EEG) (Ye et al., 2022), electrocardiogram (ECG) (Hsu et al., 2017), and electromyogram (EMG) (Cheng and Liu, 2008), to detect emotional states. While many studies have concentrated on identifying basic emotions using the valence-arousal (VA) (Zhu et al., 2023) or valence-arousal-dominance (VAD) (Lang, 1995) model, there is an increasing awareness of the complexity and mixed nature of human emotions. Research indicates that individuals may simultaneously experience multiple emotions with varying intensities at the same time (Williams and Aaker, 2002; Zhao et al., 2020). However, most existing emotion recognition methods tend to focus on single-class models (Jia et al., 2019; Zhang et al., 2018; Zheng and Lu, 2016), which may not fully capture the varied, complex, and sometimes ambiguous nature of human emotions. Accurate identification of these complex emotional states is essential, as it reflects more realistic human experiences and interactions, underscoring the need for advanced emotion recognition systems that can interpret these nuances effectively.
Figure 1. The challenges in emotion distribution learning and illustration of the proposed method. (a) cross-modal heterogeneity, (b) label correlation, (c) the schematic diagram of the proposed framework.
In contrast with multi-label emotion recognition (Li and Deng, 2019; Zhang et al., 2020, 2021a), which typically classifies emotions into distinct and singular categories such as happiness, sadness, and anger, mixed emotion recognition aims to capture the complexity of coexisting emotions that do not easily fit into traditional labels. Furthermore, mixed emotion recognition evaluates the intensity of each emotion present, whereas multi-label approaches merely identify the presence or absence of each emotional state. Jia et al. (Jia et al., 2019) adopted an emotion distribution learning method that exploits label correlations locally for facial expression recognition. Zhang et al. (Zhang et al., 2018) proposed a multi-task convolutional neural network for text emotion analysis. Existing single-modal methods struggle to fully utilize the complementary and correlative aspects between different modalities, overlooking the fact that emotions can be expressed through a combination of physiological responses and external manifestations.
Researchers have explored the efficacy of integrating multi-modal data to enhance emotion recognition performance. Peng et al. (Peng et al., 2024) proposed CARAT to model fine-grained modality-to-label dependencies and exploit the modality complementarity by a shuffle-based aggregation strategy for multi-modal multi-label emotion recognition. Shu et al. (Shu et al., 2022) proposed an EDL convolutional neural network to predict the emotion distribution from the stacked physiological signals. Liu et al. (Liu et al., 2023) developed EmotionDict to learn a set of basic emotion elements in the latent feature space for emotion distribution learning. However, as the diagram shown in Figure 1 (a), these methods ignore the intrinsic correlation and heterogeneity among different modalities, resulting in a sub-optimal recognition performance. What’s more, they cannot fully exploit the label correlation across different basic emotions. As can be seen from Figure 1 (b), “afraid” and “scared” may be highly correlated. Leveraging the label correlations as a guidance can enhance the performance of emotion distribution learning.
To address the above limitations, in this paper, we propose a multi-modal emotion distribution learning framework, named HeLo, as depicted in Figure 1 (c), which effectively mines the cross-modal heterogeneity within physiological and behavioral data, facilitating the learning of the label correlation to guide emotion distribution learning. In brief, the main contributions of our proposed model can be summarized as follows:
  • Due to differences in heterogeneity across modalities, we first adopt a cross-attention mechanism to fuse the physiological data. Then, an optimal transport (OT)-based heterogeneity mining module is devised to effectively fuse the physiological and behavioral representations.
  • For the learning of label correlation, we introduce a learnable label embedding, which is constrained by its learnable label correlation and ground-truth label correlation. Furthermore, the learnable label embeddings and label correlation are integrated through a novel label correlation-driven cross-attention mechanism.
  • Extensive experiments on two multi-modal emotion datasets demonstrate that our proposed model outperforms state-of-the-art models in both subject-dependent and subject-independent scenarios.
2. Related Work
2.1. Label Distribution Learning
In order to address the problem that both the labels and their intensities are necessarily needed, Geng et al. systematically proposed label distribution learning (Geng, 2016). According to the proposed LDL strategies, traditional LDL methods can be divided into three categories, i.e., Problem Transformation (PT) (Geng and Ji, 2013; Geng et al., 2014), Algorithm Adaptation (AA) (Geng et al., 2013, 2010), and Specialized Algorithms (SA) (Geng, 2016; Geng et al., 2013). Recently, Wen et al. proposed the learning objectives and evaluation metrics for label distribution learning, namely Cumulative Absolute Distance (CAD), Quadratic Form Distance (QFD), and Cumulative Jensen-Shannon divergence (CJS). LDL-LRR (Jia et al., 2023) adopted a novel ranking loss function for label distribution characteristics and combined it with KL-divergence as the loss term of the objective function. Gao et al. (Gao et al., 2017) proposed deep label distribution learning (DLDL), which utilizes label ambiguity in both feature learning and classifier learning to establish an end-to-end LDL framework.
To improve the performance of LDL, researchers have made attempts to exploit the label correlations. Jia et al. (Jia et al., 2019) employed a local low-rank structure to capture the label correlation implicitly to enhance facial expression recognition. Kou et al. (Kou et al., 2024) proposed TLRLDL, which leverages low-rank label correlations specifically within an auxiliary multi-label learning framework. Different from these methods, we develop an end-to-end EDL framework, which makes full use of the complementarity between physiology and behavior signals, and alleviates the negative impact of discrete emotional ambiguity on EDL through label correlation learning.
2.2. Muli-modal Emotion Recognition
As multi-modal data provide a more comprehensive understanding of human emotional states (Ngai et al., 2022; Zheng et al., 2023), multi-modal emotion recognition has emerged as an important topic in the field of affective computing. The dominant multi-modal emotion recognition methods can be divided into two subfields: single-label emotion recognition (SLER) and multi-label emotion recognition (MLER). SLER recognized emotions using a single emotion label for each discrete time. MAET (Jiang et al., 2023) processed multi-modal inputs with specialized attention modules and alleviates subject discrepancy via adversarial training. While the goal of MLER is to assign an arbitrary number of emotion category labels to an input sample. Zhang et al. (Zhang et al., 2020) integrated transformer and multi-head modality attention to effectively fuse text, visual, and audio data and predict emotions. Compared with SLER and MLER, emotion distribution learning (EDL) has gained increasing attention, as it evaluates the intensity of each basic emotion, which is crucial for a comprehensive representation of complex emotional space. EmotionDict (Liu et al., 2023) utilized behavior and physiological responses to offer complementary information for emotion recognition and disentangled emotion features of an emotional state into a weighted combination of a set of basic emotion elements. Shu et al.(Shu et al., 2022) proposed EDLConv the combine multi-modal peripheral physiological signals for emotion distribution learning. In this paper, we propose a novel EDL framework for integrating multi-modal emotional data, and the model’s performance is improved with the guidance of learned label correlation.
Figure 2. The pipeline of the proposed HeLo. (a) Modality fusion stage: We first adopt a cross-attention mechanism to fuse the physiological data. Then the OT-based heterogeneity mining module is introduced to integrate the physiological and behavioral representations. (b) Label correlation learning stage: A learnable label embedding is introduced to explore the label correlation that can be combined with the multi-modal features through a novel label correlation-driven cross-attention mechanism.
3. Methodology
The overall framework of the proposed method is illustrated in Figure 2, which consists of physiological signals fusion, OT-based heterogeneity mining, and label correlation-driven cross-attention. We will briefly introduce these components in the remainder of this section.
3.1. Multi-Modal Physiological Signals Fusion
Taking the DMER dataset as an example, after data pre-processing, we get a sampled input $x$ composed of four modalities, i.e., $x^{e}\in\mathbb{R}^{C_{e}\times d_{e}}$ for EEG, $x^{g}\in\mathbb{R}^{C_{g}\times d_{g}}$ for GSR, $x^{p}\in\mathbb{R}^{C_{p}\times d_{p}}$ for PPG, and $x^{v}\in\mathbb{R}^{C_{v}\times d_{v}}$ for video, where $C_{e}$, $C_{g}$, $C_{p}$, and $C_{v}$ denote the number of signal channels, and $d_{e}$, $d_{g}$, $d_{p}$, and $d_{v}$ are the extracted feature dimensions. As the heterogeneity within physiological signals (i.e., EEG, GSR, PPG, ECG, EDA, EMG) may be easier to handle than that of physiological and behavioral signals, we adopt a multi-head cross-attention mechanism to effectively fuse the physiological signals. Specifically, we first project EEG, GSR, and PPG signals into the same feature dimension. Then, we generate Queries on $x^{e}\in\mathbb{R}^{C\times d}$ and Keys, Values on $x^{g}\in\mathbb{R}^{C\times d}$ and $x^{p}\in\mathbb{R}^{C\times d}$, which can be formulated as follows:
(1) $\begin{gathered}Q^{e}=x^{e}\mathbf{W}_{Q}^{e},K^{m}=x^{m}\mathbf{W}_{K}^{m},V^{m}=x^{m}\mathbf{W}_{V}^{m}\end{gathered}$
where $\mathbf{W}_{Q}^{e}$, $\mathbf{W}_{K}^{m}$, $\mathbf{W}_{V}^{m}$ are learnable matrix, $m=\{g,p\}$ denotes the specific modality. Then, the cross-modal representation can be obtained by the self-attention mechanism:
(2) $x^{e\rightarrow g}=\operatorname{softmax}\left(\frac{Q^{e}\left(K^{g}\right)^{T}}{\sqrt{d}}\right)V^{g}$
$x^{e\rightarrow p}=\operatorname{softmax}\left(\frac{Q^{e}\left(K^{p}\right)^{T}}{\sqrt{d}}\right)V^{p}$
where $d$ denotes a normalization parameter equal to the dimension of $K^{m}$. To learn multiple patterns of the association within physiological signals, multi-head attention mechanism and residual connection are employed:
(3) $x_{MHA}^{e\rightarrow m}=\left(x_{h_{1}}^{e\rightarrow m}\left\|x_{h_{2}}^{e\rightarrow m}\cdots\right\|x_{h_{k}}^{e\rightarrow m}\right)\mathbf{W}_{MHA}+LN\left(x^{m}\right)$
where $x_{MHA}^{e\rightarrow m}\in\mathbb{R}^{C\times\frac{d}{h}}$ is the output of multi-head cross attention block, $h$ is the number of heads, $\mathbf{W}_{MHA}$ denotes the transformation matrix, and ||||| | represents the concatenation operation. Then, the concatenation of two cross-modal representations is adopted as the final physiological representation:
(4) $x^{phy}=\left(x_{MHA}^{e\rightarrow g}\|x_{MHA}^{e\rightarrow p}\right)$
Using EEG as the query guides the cross-attention mechanism to focus on the emotional information at the neural level, effectively capturing the relationship between EEG and other physiological signals.
3.2. Heterogeneity Mining with Optimal Transport
Physiological signals, i.e., EEG, GSR, ECG, and PPG, provide subconscious emotional responses, while behavioral data reflects discrete and context-specific insights that are influenced by conscious control and environmental interactions. It is worth noting that different data characteristics lead to significant cross-modal heterogeneity, which heightens the difficulty of multi-modal fusion. Optimal transport (OT) (Perrot et al., 2016; Cao et al., 2022; Song et al., 2024) learns the optimal matching flow between two probability distributions, which allows for the effective alignment of heterogeneous multi-modal data distributions, facilitating the extraction of coherent patterns from physiological and behavioral signals. Formally, for the previous extracted physiological representations and projected behavioral representations, $x^{Phy}\in\mathbb{R}^{2C\times d}$ and $x^{v}\in\mathbb{R}^{2C\times d}$, an optimal transport from the $x^{Phy}$ to $x^{v}$ can be defined by the discrete Kantorovich formulation (Kantorovich, 2006) to search the overall optimal matching flow:
(5) $\mathcal{WD}\left(x^{Phy},x^{v}\right)=\min_{f\in\Pi(u,\mathrm{v})}<T,\text{ Cost }>_{F}$
where $\mathcal{WD}(,)$ denotes the wasserstein distance, $\Pi\left(u,v\right)=\{T\in\mathbb{R}_{+}^{2C\times 2C}\mid T1_{2C}=u,T^{T}1_{2C}=v\}$, $1_{2C}$ represents a $2C$-dimensional all-one vector, $<,>_{F}$ refers to the Frobenius dot product, and $\text{Cost}\geq 0\in\mathbb{R}^{2C\times 2C}$ is a cost matrix calculated by the cosine distance between $x^{Phy}$ and $x^{v}$. The matrix $T\in\mathbb{R}^{2C\times 2C}$ is denoted as the optimal transport matching flow. Then, we apply the Pytorch framework-based batch-wise implementation of the Sinkhorn algorithm proposed in (Chen et al., 2020) to solve $\mathcal{WD}$. In this way, the obtained optimal transport matching flow $T$ can be used as a “cross-modal correlation” to enhance the multi-modal fusion. As such, the final multi-modal representation is adopted as follows:
(6) $x^{m}=(\text{Transformer}_{l}(x^{Phy}\otimes T)||\text{Transformer}_{l}(x^{v}))$
where $\text{Transformer}_{l}$ denotes a $l$-layer transformer encoder, $\otimes$ is the matrix multiplication, ||||| | represents the concatenation operation. The optimal transport (OT) module bridges modality heterogeneity by aligning physiological and behavioral feature spaces through optimal matching flows, thereby mitigating distribution discrepancies.
Figure 3. Illustration of the label correlation-driven cross-attention mechanism (LCDCA).
3.3. Label Correlation-driven Cross-Attention
In the emotion distribution learning scenario, correlation across different labels plays a vital role in the prediction of emotional distributions. Therefore, we design a label correlation-driven cross-attention mechanism (LCDCA), shown in Figure 3, to effectively enhance the prediction performance of the model. Specifically, we introduce a learnable label embedding $x^{L}\in\mathbb{R}^{l\times d}$, where $l$ is the number of emotion classes. To learn the label correlation, we first generate learnable label correlation $M^{L}\in\mathbb{R}^{l\times l}$ and ground-truth label correlation $M^{gt}\in\mathbb{R}^{l\times l}$ through the cosine similarity, whose $(i,j)$-th element is given as:
(7) $M_{i,j}^{L}=\frac{x_{i}^{L}\cdot x_{j}^{L}}{||x_{i}^{L}||\cdot||x_{j}^{L}||},M_{i,j}^{gt}=\frac{L_{i}\cdot L_{j}}{||L_{i}||\cdot||L_{j}||}$
where $L\in\mathbb{R}^{l\times 1}$ denotes the ground truth label distribution. Next, we adopt l-2 loss to constrain the learning of label correlation, and the correlation-constrained (CC) loss can be formulated as:
(8) $\mathcal{L}_{CC}=\|M^{L}-M^{gt}\|_{2}^{2}$
Then, we generate $Q^{L}$ on $x^{L}$, $K^{m}$ and $V^{m}$ on the projected multi-modal features $x^{m}$ through linear transformations. The label correlation-guided representation can be obtained by:
(9) $x^{o}=softmax(\frac{Q^{L}(K^{m})^{T}+M^{L}}{\sqrt{d}})V^{m}$
With the LCDCA, the proposed model can effectively utilize the label correlation for emotion distribution learning. This design allows the model to dynamically aggregate global semantic relationships between features and labels, while simultaneously incorporating sample-specific correlations through attention matrix refinement. Finally, we adopt a prediction head consisting of a three-layer MLP and a softmax layer to generate the predicted emotion distribution $\hat{L}=(d_{1},d_{2},d_{3},\cdots,d_{l})$, where $l$ is the number of emotion classes. The Kullback-Leibler divergence (KLD) is adopted as the loss function to measure the distance between the predicted emotion distribution and the ground truth emotion distribution. Therefore, the overall training of our method is:
(10) $\mathcal{L}_{Overall}=\text{KLD}(\hat{L},L)+\mathcal{L}_{CC}$
[…]
Figure 4. Predicted emotion distributions of our model and comparison methods. We show two panels of two test samples of Subject No.45. GT indicates the ground truth distribution. The numbers 1 to 10 correspond to emotions inspired, alert, excited, enthusiastic, determined, afraid, upset, nervous, scared, and distressed.
Table 1. Subject-dependent comparison of experimental results of our method and 10 comparison methods on six measures on the DMER and WESAD datasets. ↓ indicates “the smaller the better”, and ↑ indicates “the larger the better”. The parentheses show the corresponding ranks on each evaluation metric and the Average Ranks. The boldfaced scores denote the best performances and the underlined scores are the second runners. The results of the proposed HeLo are presented in a gray background.
Dataset Measure TKDE16 (Geng, 2016) TKDE16 (Geng, 2016) TPAMI13 (Geng et al., 2013) TNSRE22 (Song et al., 2022) ACMMM23 (Jiang et al., 2023) AAAI24 (Peng et al., 2024) TKDE23 (Jia et al., 2023) IJCAI24 (Kou et al., 2024) ICCV23 (Wen et al., 2023) TAFFC23 (Liu et al., 2023) Ours
PT-SVM AA-KNN SA-CPNN Conformer MAET CARAT LDL-LRR TLRLDL CAD EmotionDict
DMER Chebyshev (↓) 0.0851 (7) 0.0949 (10) 0.0917 (8) 0.0661 (3) 0.0522 (2) 0.0752 (4) 0.0936 (9) 0.0847 (6) 0.1017 (11) 0.0753 (5) 0.0446 (1)
Clark (↓) 0.5787 (6) 0.6649 (10) 0.6333 (8) 0.4743 (3) 0.3762 (2) 0.5951 (7) 0.6529 (9) 0.5380 (4) 0.7517 (11) 0.5463 (5) 0.3256 (1)
Canberra (↓) 1.5294 (7) 1.2357 (3) 1.8092 (9) 1.2581 (5) 0.9875 (2) 1.5965 (8) 1.8224 (10) 1.3271 (5) 1.9415 (11) 1.4771 (6) 0.8664 (1)
KL (↓) 0.0837 (6) 0.1409 (10) 0.1048 (8) 0.0593 (3) 0.0288 (1) 0.1137 (9) 0.0644 (4) 0.0925 (7) 0.1460 (11) 0.0766 (5) 0.0323 (2)
Cosine (↑) 0.9237 (8) 0.9315 (6) 0.9058 (9) 0.9464 (3) 0.9688 (2) 0.9332 (5) 0.9016 (10) 0.9339 (4) 0.8923 (11) 0.9313 (7) 0.9714 (1)
Intersection (↑) 0.8417 (8) 0.8425 (7) 0.8150 (9) 0.8716 (3) 0.9113 (2) 0.8474 (6) 0.8136 (10) 0.8537 (5) 0.8077 (11) 0.8492 (5) 0.9128 (1)
Average Rank (↓) 7 (7) 7.66 (8) 8.5 (9) 3.33 (3) 1.83 (2) 6.16 (6) 9.5 (10) 4.66 (4) 11 (11) 5.5 (5) 1.16 (1)
WESAD Chebyshev (↓) 0.0515 (8) 0.0346 (4) 0.0998 (11) 0.0209 (3) 0.0347 (5) 0.0354 (6) 0.0357 (7) 0.0622 (10) 0.0569 (9) 0.0090 (2) 0.0073 (1)
Clark (↓) 0.3940 (8) 0.2779 (4) 0.7285 (11) 0.1838 (3) 0.3503 (7) 0.2812 (5) 0.2822 (6) 0.5381 (10) 0.4618 (9) 0.0733 (2) 0.0653 (1)
Canberra (↓) 1.0249 (8) 0.7381 (4) 2.0746 (11) 0.4810 (3) 0.9100 (7) 0.7430 (5) 0.7466 (6) 1.3586 (10) 1.2199 (9) 0.1876 (2) 0.1614 (1)
KL (↓) 0.0320 (8) 0.0243 (5) 0.1226 (10) 0.0110 (3) 0.0251 (6) 0.0210 (4) 0.0273 (7) 0.2051 (11) 0.0440 (9) 0.0017 (2) 0.0010 (1)
Cosine (↑) 0.9728 (8) 0.9793 (7) 0.8918 (11) 0.9909 (3) 0.9867 (4) 0.9824 (5) 0.9821 (6) 0.9541 (10) 0.9624 (9) 0.9985 (2) 0.9992 (1)
Intersection (↑) 0.9024 (8) 0.9297 (4) 0.7896 (11) 0.9550 (3) 0.9213 (7) 0.9296 (5) 0.9292 (6) 0.8746 (10) 0.8812 (9) 0.9820 (2) 0.9905 (1)
Average Rank (↓) 8 (8) 5.33 (5) 10.83 (11) 3 (3) 6 (6) 5 (4) 6.33 (7) 10.16 (10) 9 (9) 2.5 (2) 1 (1)
Table 2. Subject-independent comparison of experimental results of our method and 10 comparison methods on six measures on the DMER and WESAD datasets. ↓ indicates “the smaller the better”, and ↑ indicates “the larger the better”. The parentheses show the corresponding ranks on each evaluation metric and the Average Ranks. The boldfaced scores denote the best performances and the underlined scores are the second runners. The results of the proposed HeLo are presented in a gray background.
Dataset Measure TKDE16 (Geng, 2016) TKDE16 (Geng, 2016) TPAMI13 (Geng et al., 2013) TNSRE22 (Song et al., 2022) ACMMM23 (Jiang et al., 2023) AAAI24 (Peng et al., 2024) TKDE23 (Jia et al., 2023) IJCAI24 (Kou et al., 2024) ICCV23 (Wen et al., 2023) TAFFC23 (Liu et al., 2023) Ours
PT-SVM AA-KNN SA-CPNN Conformer MAET CARAT LDL-LRR TLRLDL CAD EmotionDict
DMER Chebyshev (↓) 0.1462 (11) 0.1070 (10) 0.0928 (3) 0.0933 (4) 0.0939 (5) 0.1023 (9) 0.1012 (8) 0.0964 (6) 0.0925 (2) 0.0945 (7) 0.0882 (1)
Clark (↓) 1.2877 (11) 0.6860 (9) 0.6440 (3) 0.6515 (4) 0.6605 (5) 0.6909 (10) 0.6856 (8) 0.6702 (6) 0.6379 (2) 0.6676 (7) 0.6289 (1)
Canberra (↓) 2.0459 (11) 1.8308 (7) 1.8542 (8) 1.7689 (2) 1.7863 (3) 1.9161 (10) 1.9071 (9) 1.8244 (6) 1.8104 (5) 1.8092 (4) 1.7603 (1)
KL (↓) 0.5528 (11) 0.1227 (9) 0.1089 (5) 0.1061 (2) 0.1088 (4) 0.1245 (10) 0.1222 (8) 0.1142 (7) 0.1076 (3) 0.1104 (6) 0.1027 (1)
Cosine (↑) 0.8295 (11) 0.8843 (10) 0.9022 (4) 0.9038 (2) 0.9016 (5) 0.8899 (9) 0.8917 (8) 0.8964 (7) 0.9036 (3) 0.9001 (6) 0.9148 (1)
Intersection (↑) 0.7762 (11) 0.8050 (8) 0.8102 (6) 0.8161 (2) 0.8147 (3) 0.8026 (10) 0.8037 (9) 0.8087 (7) 0.8122 (4) 0.8119 (5) 0.8194 (1)
Average Rank (↓) 11 (11) 8.83 (9) 4.83 (5) 2.66 (2) 4.16 (4) 9.66 (10) 8.33 (8) 6.5 (7) 3.16 (3) 5.83 (6) 1 (1)
WESAD Chebyshev (↓) 0.0471 (7) 0.0500 (9) 0.0894 (11) 0.0436 (3) 0.0516 (10) 0.0454 (5) 0.0466 (6) 0.0441 (4) 0.0481 (8) 0.0421 (2) 0.0403 (1)
Clark (↓) 0.4506 (8) 0.4516 (10) 0.6394 (11) 0.4048 (3) 0.4514 (9) 0.4270 (5) 0.4471 (6) 0.4145 (4) 0.4488 (7) 0.3631 (2) 0.3455 (1)
Canberra (↓) 1.1928 (10) 1.1762 (7) 1.7384 (11) 1.0806 (4) 1.1666 (6) 1.1874 (9) 1.1817 (8) 1.0746 (3) 1.1647 (5) 0.9530 (2) 0.9329 (1)
KL (↓) 0.0347 (5) 0.0423 (9) 0.0944 (11) 0.0369 (7) 0.0445 (10) 0.0345 (4) 0.0343 (3) 0.0372 (8) 0.0366 (6) 0.0282 (1) 0.0283 (2)
Cosine (↑) 0.9727 (5) 0.9663 (10) 0.9154 (11) 0.9699 (7) 0.9678 (9) 0.9731 (3) 0.9731 (3) 0.9695 (8) 0.9711 (6) 0.9773 (2) 0.9790 (1)
Intersection (↑) 0.8935 (7) 0.8933 (9) 0.8254 (11) 0.9015 (3) 0.8967 (4) 0.8944 (6) 0.8945 (5) 0.8917 (10) 0.8935 (7) 0.9126 (2) 0.9154 (1)
Average Rank (↓) 7 (8) 9 (10) 11 (11) 4.5 (3) 8 (9) 5.33 (5) 5.16 (4) 6.16 (6) 6.5 (7) 1.83 (2) 1.16 (1)
4. Experiment
4.1. Multi-Modal Emotion Datasets
We evaluate our proposed method on two publicly available datasets: DMER (Yang et al., 2024) and WESAD (Schmidt et al., 2018) datasets. DMER dataset contains four modalities including EEG, galvanic skin response (GSR), photoplethysmogram (PPG), and facial videos collected from 80 participants. Each participant was asked to watch 32 video clips. In each trial, the subject first watched one video clip, and then completed the 10-item short positive and negative affect schedules (PANAS (Watson, 1994)). The score for each basic emotion ranged from 1 (none) to 5 (strong) and was then transformed into emotion distributions. The data of 73 subjects were used for further experiments in this paper. WESAD contains ECG, EMG, electrodermal activity (EDA), and 3-axis accelerometer (ACC) from 17 participants. The data were collected at a sampling rate of 700Hz, and the goal was to study four different affective states namely neutral, stressed, amused, and meditated. Upon completion of each trial, the ground truth labels for the affect states were collected using the PANAS (Watson, 1994). 14 participants were utilized in our experiments. Notably, ECG, EMG, and EDA are the ”Physiological data”, and ACC denotes the ”Behavioral Data” in our proposed method. Details of the data preprocessing can be referred to the supplementary materials.
4.2. Evaluation Metrics
For the evaluation of our proposed model, following the previous emotion distribution learning studies (Shu et al., 2022; Liu et al., 2023; Geng, 2016), we adopt six distribution-based measurements, including four distance metrics (Chebyshev (↓), Clark (↓), Canberra (↓), and Kullback-Leibler (KL) (↓)) and two similarity measurements (Cosine (↑) and Intersection (↑)). ↓ indicates the metric is “the smaller the better”, and ↑ denotes “the larger the better”. Moreover, we reported the Average Rank denoting the mean rank of the six metrics following (Shu et al., 2022; Liu et al., 2023).
4.3. Implementation Details
In our experiments, following previous works (Shu et al., 2022; Liu et al., 2023), both subject-dependent and subject-independent settings are adopted to verify the method on the two datasets. In subject-dependent recognition, the data of each subject are split into the training set and testing set in a ratio of 8:2 randomly, and the average performance across all subjects is taken as the final result. For subject-independent recognition, we utilize the leave-one-subject-out cross-validation for the evaluation, and the average performance across all iterations is reported as the final result. Our proposed framework is trained on a 12GB NVIDIA GeForce RTX3080 GPU and developed under PyTorch. The training adopts Adam optimizer with a learning rate of $10^{-3}$. The batch size is set to 128, the number of epochs for training is set to 300, and the number of attention heads is 4. The embedding size in the attention mechanism and feed-forward network is set to 128, and 64, respectively, and we set the depth of transformer blocks to 1. Our code will be released soon.
4.4. Comparison with State-of-the-art Methods
To evaluate the effectiveness of our proposed method, we compare our model with several state-of-the-art models on both the DMER and AMIGOS datasets. These methods including label distribution learning methods (PT-SVM (Geng, 2016), AA-KNN (Geng, 2016), SA-CPNN (Geng et al., 2013)), EEG-based classification method (Conformer (Song et al., 2022)), multi-modal emotion recognition method (MAET (Jiang et al., 2023)), multi-modal multi-label emotion recognition method (CARAT (Peng et al., 2024)), recently proposed state-of-the-art single-modal label distribution learning methods (LDL-LRR (Jia et al., 2023), TLRLDL (Kou et al., 2024), CAD (Wen et al., 2023)), and multi-modal emotion distribution learning method (EmotionDict (Liu et al., 2023)).
The results of the subject-dependent recognition on the DMER and WESAD datasets are summarized in Table 1. As can be seen from the table, except for the KL metric on the DMER dataset, our proposed method outperforms all the other state-of-the-art methods over the six metrics on both datasets. Benefiting from deep learning’s ability to extract high-dimensional features, the deep learning-based methods achieve relatively better performance than the traditional LDL methods (PT-SVM, AA-KNN, and SA-CPNN). Besides, visualization examples of the predicted emotion distributions using all the methods and the ground truths on the DMER dataset are shown in Figure 4. It can be seen that some methods, i.e., PT-SVM, SA-CPNN, Conformer, and MAET, are more likely to output similar distributions for different signal samples, leading to sub-optimal emotion distribution learning results. The qualitative results show that our method can predict emotion distributions most similar to the ground truths. The superiority of our method lies in that our model not only effectively utilizes the fusion of heterogeneous multi-modal data, but also leverages the learnable label correlation for the prediction of emotion distribution.
Table 3. Ablation study in modalities on the DMER.
Settings Measure w/o EEG w/o GSR w/o PPG w/o Video Ours
Subject-Dependent Chebyshev (↓) 0.0529 0.0574 0.0487 0.0552 0.0446
Clark (↓) 0.3372 0.3550 0.3671 0.3380 0.3256
Canberra (↓) 0.9028 0.8987 0.8820 0.9065 0.8664
KL (↓) 0.0381 0.0429 0.0357 0.0377 0.0323
Cosine (↑) 0.9677 0.9591 0.9692 0.9628 0.9714
Intersection (↑) 0.9102 0.9007 0.9086 0.9033 0.9128
Subject-Independent Chebyshev (↓) 0.0977 0.0953 0.0964 0.0926 0.0882
Clark (↓) 0.6906 0.6742 0.6470 0.6635 0.6289
Canberra (↓) 1.9261 1.9077 1.8074 1.8250 1.7603
KL (↓) 0.1175 0.1244 0.1085 0.1134 0.1027
Cosine (↑) 0.9080 0.9075 0.9122 0.9107 0.9148
Intersection (↑) 0.8197 0.8056 0.8138 0.8121 0.8194
Table 4. Ablation study in modalities on the WESAD.
Settings Measure w/o ECG w/o EMG w/o EDA w/o ACC Ours
Subject-Dependent Chebyshev (↓) 0.0088 0.0074 0.0079 0.0082 0.0073
Clark (↓) 0.0744 0.0722 0.0762 0.0695 0.0653
Canberra (↓) 0.1730 0.1630 0.1658 0.1725 0.1614
KL (↓) 0.0022 0.0016 0.0014 0.0016 0.0010
Cosine (↑) 0.9915 0.9960 0.9955 0.9971 0.9992
Intersection (↑) 0.9873 0.9899 0.9892 0.9885 0.9905
Subject-Independent Chebyshev (↓) 0.0502 0.0419 0.0429 0.0474 0.0403
Clark (↓) 0.3877 0.3597 0.3721 0.3842 0.3455
Canberra (↓) 1.0748 0.9516 0.9579 1.0029 0.9329
KL (↓) 0.0322 0.0308 0.0344 0.0315 0.0283
Cosine (↑) 0.9688 0.9722 0.9641 0.9658 0.9790
Intersection (↑) 0.8870 0.9095 0.9017 0.8991 0.9154
Table 2 illustrates the subject-independent evaluation of the comparison methods and our proposed model on the DMER and WESAD datasets. Except for the KL metric on the WESAD dataset, our method achieves the best results across both datasets and shows the highest average rank. The subject-independent recognition is more challenging than the subject-dependent situation, showing the superiority of our proposed method in capturing the subject-invariant emotional features across different individuals.
(a) DMER Subject-Dependent
(b) DMER Subject-Indpendent
(c) WESAD Subject-Dependent
(d) WESAD Subject-Independent
Figure 5. Ablation of network components in the proposed method. The experiment is conducted on the DMER and WESAD datasets, using both subject-dependent and subject-independent evaluation.
4.5. Ablation Studies
Ablation on network components. To verify the rationality of the modules of the proposed framework, we conduct ablation experiments on the DMER and WESAD datasets by removing each module of our method, i.e., cross-attention-based physiological signals fusion (CAPF), optimal transport-based heterogeneity mining (OTHM), and label-correlation-driven cross-attention (LCDCA), and the results are illustrated in Figure 5. The experimental results show that each module in the model has a positive contribution to the emotion distribution learning performance of the model, and the proposed method achieves the best performance through the effective integration of each module. CAPF effectively fuses the physiological signals to enhance the model’s performance. OTHM achieves better performance than concatenation, indicating its proficiency in capturing heterogeneous information across physiological and behavioral features. Furthermore, LCDCA greatly improves the model’s performance as it integrates the label correlations for the learning of emotion distribution.
Ablation on the modalities. We also assessed the impact of each modality by separately removing each modality from our full model, the results are shown in Table 3 and Table 4. These experiments were conducted in both subject-dependent and subject-independent settings across both datasets. The results indicate that each modality contributes significantly to the overall performance of multi-modal emotion distribution learning, effectively integrating multiple modalities can enhance the emotion distribution performance. More ablation results on modalities can be referred to the supplementary materials.
Figure 6. Effect of the number of attention heads. It shows the performance over six metrics on the DMER dataset.
Figure 7. The effect of the depth of transformer block in the model on six evaluation metrics on the DMER dataset.
Ablation on the number of attention heads. The number of attention heads is usually an essential parameter for models based on attention mechanisms. Thus, we also compared the effect of different choices of the number of heads on the six evaluation metrics of the model on the DMER dataset in both subject-dependent and subject-independent settings, shown in Figure 6. It can be seen that the model’s performance first increases across subject-dependent and subject-independent settings as the number of attention heads increases. Then, the model achieves the best performance across all six evaluation metrics when the number of attention heads is equal to 4, after which the performance declines as the number of heads continues to increase. However, the number of attention heads has a relatively limited impact on the performance of the model.
Ablation on the depth of transformer block. The depth of end-to-end models is considered a crucial factor in determining their fitting ability. Therefore, we evaluate the model while varying the depth of the block on the DMER dataset to explore the impact of the depth of the transformer block. Figure 7 shows the optimal classification performance was achieved when the depth of the block was 1, and as the depth increased, the classification accuracy tended to slightly decrease and leveled off as the depth continued to increase, as the greater the depth of the transformer block, the more likely it is to overfit the model.
Figure 8. Visualization of the feature embedding.
Figure 9. Label correlations visualization. A higher red intensity value indicates a stronger correlation.
4.6. Visualization
Visualization of the cross-modal representations. We visualize the multi-modal feature space in the testing set by utilizing t-SNE (Van der Maaten and Hinton, 2008), and the results are illustrated in Figure 8. In Figure 8 (a), before the OTHM module, there is a noticeable separation between physiological (blue) and behavioral (red) representations. This pronounced heterogeneity illustrates the challenges faced in multi-modal fusion due to distinct feature distributions, which can impede effective integration and analysis. With the help of OTHM, shown in Figure 8 (b), the cross-modal representations exhibit a closer arrangement and overlap more significantly than before. This reduction in heterogeneity suggests that OTHM has effectively minimized the disparities between the physiological and behavioral features, facilitating a better-aligned feature space and a more coherent multi-modal fusion.
Visualization of the learned label correlations. To verify the effectiveness of the proposed LCDCA module, we visualize the learned label correlation $M^{L}$ in Figure 7 to illustrate the interpretability, and the results on the testing set of DMER dataset are shown in Figure 9. As can be seen from Figure 9, the label correlations differ from different trials and different subjects, which jointly attend to rich semantic information from various perspectives. From the horizontal view, “afraid” is highly correlated with “nervous”, “scared”. Besides, “enthusiastic”, and “determined” are usually correlated with “excited”, “alert”. All of these conclusions are consistent with human intuition, so as to demonstrate the better emotion distribution learning performance of our proposed model.
Table 5. Comparison between our method and other models used in our experiments in terms of network efficiency.
Metric TNSRE22 (Song et al., 2022) ACMMM23 (Jiang et al., 2023) AAAI24 (Peng et al., 2024) TAFFC23 (Liu et al., 2023) Ours
Conformer MAET CARAT EmotionDict
FLOPs 2.422M 91.955M 330.671M 156.354M 11.656M
Params 1.023M 15.672M 6.780M 0.483M 4.040M
4.7. Model Efficiency
We compared the efficiency of our model with 10 other methods. On one hand, deep learning-based methods generally achieve better emotion distribution learning performance than traditional label distribution learning methods. On the other hand, traditional label distribution learning typically have fewer parameters. Therefore, we conducted a comparison of FLOPs (Floating Point Operations per Second) and the number of parameters that need to be learned in the model with the other four deep learning-based methods, i.e., Conformer (Song et al., 2022), MAET (Jiang et al., 2023), CARAT (Peng et al., 2024), and EmotionDict (Liu et al., 2023)). The results are summarized in Table 5.
5. Conclusion
In this paper, we introduce a novel multi-modal emotion distribution learning framework, HeLo, which effectively addresses the challenges of capturing complex, mixed emotional states from heterogeneous physiological and behavioral data. The CAPF module ensures the robust fusion of multi-modal physiological signals, and the OTHM module efficiently mines the heterogeneity between physiological and behavioral representations, bridging the gap between these diverse data sources and enhancing feature alignment. Moreover, the LCDCA fully exploits label correlations, guiding the learning of emotion distributions in a semantically meaningful way. Experimental results on two multi-modal emotion datasets demonstrate the efficacy of our proposed method. In the future, we plan to utilize LLMs to extract the label embeddings, which provide strong semantic priors to our model instead of random initialization.
Supplementary Material
Appendix A Dataset Preprocessing
A.1. DMER Dataset
In the DMER dataset, we follow the preprocessing pipelines in the original paper (Yang et al., 2024). For EEG signals, differential entropy (DE) features in each EEG channel from 5 frequency bands ($\delta$ band (1-3Hz), $\theta$ band (4-7Hz), $\alpha$ band (8-13Hz), $\beta$ band (14-30Hz) and $\gamma$ band (31-50Hz)) were extracted. For GSR signals, the median, mean, standard deviation, minimum, maximum, ratio of minimum, and the ratio of maximum of the raw GSR signal and its first-order and second-order derivatives were extracted as time domain statistical features according to the feature extraction (Udovičić et al., 2017). Then, the Fast Fourier Transform (FFT) was adopted to transform the GSR signals from the time domain into the frequency domain, and the median, mean, standard deviation, maximum, minimum, and range of the signal were extracted (Udovičić et al., 2017). In addition, the power spectral density (PSD) of the frequency band [0, 2] Hz was extracted using Welch’s power spectral density. The same features except PSD were extracted for the PPG signals. Notably, the size of the sliding window for both GSR and PPG feature extraction was 5 seconds with an 80% overlap between two consecutive windows (Zhang et al., 2021b). For facial video data, local binary patterns from three orthogonal planes (LBP-TOP) (Zhao and Pietikainen, 2007; Yan et al., 2014) were used for the feature extraction. Due to the missing data in subject 3, 4, 13, 16, 17, 27, and 31, the data of the remaining 73 subjects are chosen for further experiments. After all the preprocessing steps, a 913-dimension feature vector, including 90 EEG features, 28 GSR features, 27 PPG features, and 768 video features for a specific subject was obtained for further experiments. The summary of extracted features is listed in Table 6.
A.2. WESAD Dataset
In the WESAD dataset, features were computed with a window size of five seconds. The feature extraction methods refer to (Koelstra et al., 2011; Schmidt et al., 2018). For ACC data, the mean, and standard deviation for each axis were separately and summed over all axes, absolute integral for each/all axis, and peak frequency for each axis were extracted. Following (Kim and André, 2008), the heartbeats were accurately localized in ECG signals (R-peaks) to calculate the inter beat intervals (IBI). Using IBI values, the heart rate (HR) and heart rate variability (HRV) time series were calculated. For EDA data, the mean and standard deviation of the signal, the number of skin conductance response (SCR) peaks, and the mean amplitude of these peaks were extracted. For EMG data, the mean, standard deviation, range, integral, and median of the signal were extracted, along with the 10th and 90th percentiles. In the frequency domain, the peak frequency and total power spectral density were computed. Additionally, the number of peaks, mean peak amplitude, the standard deviation of peak amplitudes, the sum of peak amplitudes, and the normalized sum of peak amplitudes were calculated. Due to sensor malfunction in subject 1, and 12 and missing labels in subject 5, the data of the remaining 14 participants was used for experiments. After all the preprocessing steps, we obtained 12 ACC features, 73 ECG features, 4 EDA features, and 14 EMG features. The summary of extracted features is illustrated in Table 7.
Table 6. Extracted Affective Features for each Modality (feature dimension stated in parenthesis) on the DMER dataset.
Modality Extracted features
EEG (90)
5 bands (sigma, theta, alpha, beta, gamma) differential
entropy (DE) features for each EEG channel.
GSR (28)
Time domain: median, mean, standard deviation, minimum,
maximum, ratio of minimum, ratio of maximum of the raw
GSR signal, first-order and second-order derivatives.
Frequency domain: fast fourier transform (FFT), median,
mean, standard deviation, maximum, minimum, range of
the signal, PSD of frequency band [0, 2] Hz.
PPG (27)
Time domain: median, mean, standard deviation, minimum,
maximum, ratio of minimum, ratio of maximum of the raw
GSR signal, first-order and second-order derivatives.
Frequency domain: fast fourier transform (FFT), median,
mean, standard deviation, maximum, minimum, range of
the signal.
Video (768)
Local binary patterns from three orthogonal planes (LBP-TOP)
features from each video frame.
Table 7. Extracted Affective Features for each Modality (feature dimension stated in parenthesis) on the WESAD dataset.
Modality Extracted features
ECG (73)
Inter beat intervals (IBI), heart rate (HR), and heart rate
variability (HRV) time series.
EDA (4)
mean/standard deviation of the signal, the number of skin
conductance response (SCR) peaks, the mean amplitude
of SCR peaks.
EMG (14)
mean, standard deviation, range, integral, median of the
signal along with the 10th and 90th percentiles, peak
frequency and total power spectral density, the number of
peaks, mean peak amplitude, standard deviation of peak
amplitudes, sum of peak amplitudes, normalized sum of
peak amplitudes were calculated.
ACC (12)
standard deviation for each axis separately and summed
over all axes, absolute integral for each/all axis, peak
frequency for each axis.
Appendix B Details of the Comparison Methods
  • PT-SVM (Geng, 2016) are based on Bayes and SVM classifiers with problem transformation strategy, which converts LDL to single-label distribution problems via transforming the training data into weighted single-label data.
  • AA-KNN (Geng, 2016) is an algorithm adaptation strategy-based method, which extend classic k-Nearest Keighbor (KNN) to handle LDL problems.
  • SA-CPNN (Geng et al., 2013) is a specialized algorithm for LDL, which is built on a three-layer conditional probability neural network.
  • Conformer (Song et al., 2022) adopts convolutional neural networks (CNNs) and Transformer architectures to mine local and global dependencies in sequential data, respectively.
  • MAET (Jiang et al., 2023) utilizes specialized modules for processing uni-modal or multi-modal inputs flexibly, and the subject discrepancy is alleviated by adopting adversarial training.
  • CARAT (Peng et al., 2024) adopts a reconstruction-based fusion mechanism to better model fine-grained modality-to-label dependencies by contrastively learning modal-separated and label-specific features. To further exploit the modality complementarity, a shuffle-based aggregation strategy was introduced to enrich co-occurrence collaboration among labels.
  • LDL-LRR (Jia et al., 2023) adopts a novel ranking loss function for label distribution characteristics, and combines it with KL-divergence as the loss term of the objective function.
  • TLRLDL (Kou et al., 2024) focuses on improving label distribution learning (LDL) by leveraging low-rank label correlations specifically within an auxiliary multi-label learning framework.
  • CAD (Wen et al., 2023) computes the minimal cost required to transform one label distribution into another by considering these spatial relationships label-by-label.
  • EmotionDict (Liu et al., 2023) adopts modality attention to learn cross-modal interactions and disentangle the mixed emotion representations into a weighted sum of a set of basic emotion elements in a shared latent space to enhance the emotion distribution learning performance.
Notably, the label distribution learning methods, i.e., LDL-LRR (Jia et al., 2023), TLRLDL (Kou et al., 2024), CAD (Wen et al., 2023), is implemented using the PyLDL toolkit in our experiments.
Table 8. Ablation of Query Modality on the DMER dataset
Measure GSR PPG EEG
Cheby.-dependent (↓) 0.0574 0.0591 0.0446
Cl.-dependent (↓) 0.3882 0.4045 0.3256
Canb.-dependent (↓) 0.9219 0.9182 0.8664
KL.-dependent (↓) 0.0446 0.0462 0.0323
Cos.-dependent (↑) 0.9590 0.9498 0.9714
Inter.-dependent (↑) 0.8812 0.8990 0.9128
Cheby.-independent (↓) 0.0942 0.0924 0.0882
Cl.-independent (↓) 0.7210 0.6970 0.6289
Canb.-independent (↓) 1.8473 1.8821 1.7603
KL.-independent (↓) 0.1183 0.1196 0.1027
Cos.-independent (↑) 0.9055 0.9005 0.9148
Inter.-independent (↑) 0.7992 0.8026 0.8194
Appendix C Analysis on Query Modality
Compared with peripheral physiological signals such as GSR and PPG, EEG is more closely related to the neural mechanism of emotion generation. Using EEG as the query guides the cross-attention mechanism to focus on the emotional information at the neural level, effectively capturing the relationship between EEG and other physiological signals. We conduct ablation experiments of query modality on the DMER dataset with both subject-dependent and subject-independent settings. The ablation results were shown in Table 8, demonstrating EEG’s efficacy and appropriateness for this role in our method’s architecture.
Table 9. Ablation study in modalities of subject-dependent experiment on the DMER dataset.
Modality Measure
EEG GSR PPG Video Cheby. (↓) Cl. (↓) Canb. (↓) KL. (↓) Cos. (↑) Inter. (↑)
0.0591 0.3722 0.8825 0.0442 0.9407 0.8896
0.0695 0.4015 0.9210 0.0584 0.9318 0.8685
0.0688 0.4128 0.9256 0.0559 0.9261 0.8602
0.0602 0.3885 0.9194 0.0521 0.9335 0.8725
0.0571 0.3689 0.9112 0.0431 0.9429 0.8942
0.0580 0.3660 0.9088 0.0472 0.9446 0.8921
0.0548 0.3576 0.9011 0.0409 0.9509 0.8995
0.0595 0.3915 0.9117 0.0518 0.9445 0.8821
0.0628 0.3874 0.9050 0.0496 0.9420 0.8870
0.0609 0.4042 0.9214 0.0487 0.9392 0.8804
0.0552 0.3380 0.9065 0.0377 0.9628 0.9086
0.0574 0.3550 0.8987 0.0429 0.9591 0.9007
0.0487 0.3671 0.8820 0.0357 0.9692 0.9086
0.0529 0.3372 0.9028 0.0381 0.9677 0.9102
0.0446 0.3256 0.8664 0.0323 0.9714 0.9128
Table 10. Ablation study in modalities of subject-independent experiment on the DMER dataset.
Modality Measure
EEG GSR PPG Video Cheby. (↓) Cl. (↓) Canb. (↓) KL. (↓) Cos. (↑) Inter. (↑)
0.0991 0.6998 1.9258 0.1205 0.8901 0.7962
0.1126 0.7218 1.9662 0.1309 0.8795 0.7889
0.1039 0.7150 1.9581 0.1250 0.8850 0.7835
0.1021 0.7246 1.9440 0.1276 0.8872 0.7991
0.0984 0.6762 1.8709 0.1198 0.9021 0.8083
0.0977 0.6881 1.8652 0.1186 0.8984 0.8144
0.0968 0.7017 1.8845 0.1178 0.8917 0.8098
0.1004 0.7008 1.9308 0.1210 0.8958 0.8022
0.0974 0.6925 1.9412 0.1195 0.8921 0.8035
0.0985 0.6977 1.9350 0.1226 0.8994 0.8020
0.0926 0.6635 1.8250 0.1134 0.9107 0.8121
0.0953 0.6742 1.9077 0.1244 0.9075 0.8056
0.0964 0.6470 1.8074 0.1085 0.9122 0.8138
0.0977 0.6906 1.9261 0.1175 0.9080 0.8197
0.0882 0.6289 1.7603 0.1027 0.9148 0.8194
Table 11. Ablation study in modalities of subject-dependent experiment on the WESAD dataset.
Modality Measure
ECG EMG EDA ACC Cheby. (↓) Cl. (↓) Canb. (↓) KL. (↓) Cos. (↑) Inter. (↑)
0.0122 0.0826 0.1729 0.0028 0.9720 0.9602
0.0146 0.0885 0.1890 0.0042 0.9651 0.9548
0.0137 0.0923 0.1921 0.0055 0.9622 0.9496
0.0141 0.0848 0.1847 0.0048 0.9585 0.9447
0.0106 0.0783 0.1685 0.0023 0.9876 0.9725
0.0119 0.0790 0.1699 0.0019 0.9864 0.9748
0.0097 0.0744 0.1671 0.0026 0.9909 0.9826
0.0120 0.0831 0.1772 0.0034 0.9815 0.9705
0.0128 0.0925 0.1806 0.0035 0.9840 0.9720
0.0104 0.0887 0.1825 0.0040 0.9881 0.9736
0.0082 0.0695 0.1725 0.0016 0.9971 0.9885
0.0074 0.0722 0.1630 0.0016 0.9960 0.9899
0.0079 0.0762 0.1658 0.0014 0.9955 0.9892
0.0088 0.0744 0.1730 0.0022 0.9915 0.9873
0.0073 0.0653 0.1614 0.0010 0.9992 0.9905
Table 12. Ablation study in modalities of subject-independent experiment on the WESAD dataset.
Modality Measure
ECG EMG EDA ACC Cheby. (↓) Cl. (↓) Canb. (↓) KL. (↓) Cos. (↑) Inter. (↑)
0.0528 0.3854 1.0525 0.0389 0.9603 0.8962
0.0539 0.3927 1.1884 0.0448 0.9518 0.8890
0.0548 0.4018 1.1123 0.0394 0.9422 0.8748
0.0541 0.3890 1.1091 0.0406 0.9475 0.8851
0.0520 0.3578 0.9574 0.0382 0.9642 0.9057
0.0517 0.3660 0.9971 0.0415 0.9733 0.9117
0.0510 0.3626 1.0604 0.0337 0.9715 0.9084
0.0537 0.3795 1.1072 0.0390 0.9670 0.8915
0.0518 0.3864 1.0875 0.0428 0.9506 0.8928
0.0525 0.3815 1.0256 0.0375 0.9573 0.9055
0.0474 0.3842 1.0029 0.0315 0.9658 0.8991
0.0419 0.3597 0.9516 0.0308 0.9722 0.9095
0.0429 0.3721 0.9579 0.0344 0.9641 0.9017
0.0502 0.3877 1.0748 0.0322 0.9688 0.8870
0.0403 0.3455 0.9329 0.0283 0.9790 0.9154
Appendix D More Ablation Results on Modalities
We conducted thorough ablation experiments on different combinations of different emotion modalities. The results are illustrated in Table 9, 10, 11, and 12. The experimental results demonstrate the necessity and effectiveness of our proposed multi-modal emotion distribution learning framework.

### v3

@@
-HeLo:::: Heterogeneous Multi-Modal Fusion with Label Correlation for Emotion Distribution Learning
+HeLo::: Heterogeneous Multi-Modal Fusion with Label Correlation for Emotion Distribution Learning
 Abstract.
@@
 Researchers have investigated various behavioral signals, including facial expressions (Verma et al., 2021) and voice signals (Yi and Mak, 2020), as well as physiological signals like electroencephalogram (EEG) (Ye et al., 2022), electrocardiogram (ECG) (Hsu et al., 2017), and electromyogram (EMG) (Cheng and Liu, 2008), to detect emotional states.
-While many studies have concentrated on identifying basic emotions using the valence-arousal (VA) (Zhu et al., 2023) or valence-arousal-dominance (VAD) (Lang, 1995) model, there is an increasing awareness of the complexity and mixed nature of human emotions.
+While many studies have concentrated on identifying basic emotions using the valence-arousal (VA) (Zhu et al., 2024) or valence-arousal-dominance (VAD) (Lang, 1995) model, there is an increasing awareness of the complexity and mixed nature of human emotions.
 Research indicates that individuals may simultaneously experience multiple emotions with varying intensities at the same time (Williams and Aaker, 2002; Zhao et al., 2020).
@@
 Existing single-modal methods struggle to fully utilize the complementary and correlative aspects between different modalities, overlooking the fact that emotions can be expressed through a combination of physiological responses and external manifestations.
-Researchers have explored the efficacy of integrating multi-modal data to enhance emotion recognition performance.
+Researchers have explored the efficacy of integrating multi-modal data to enhance emotion recognition performance (Liu et al., 2025).
 Peng et al.
@@
 2.2. Muli-modal Emotion Recognition
-As multi-modal data provide a more comprehensive understanding of human emotional states (Ngai et al., 2022; Zheng et al., 2023), multi-modal emotion recognition has emerged as an important topic in the field of affective computing.
+As multi-modal data provide a more comprehensive understanding of human emotional states (Ngai et al., 2022; Zheng et al., 2023; Zhu et al., 2025), multi-modal emotion recognition has emerged as an important topic in the field of affective computing.
 The dominant multi-modal emotion recognition methods can be divided into two subfields: single-label emotion recognition (SLER) and multi-label emotion recognition (MLER).
@@
 (5) $\mathcal{WD}\left(x^{Phy},x^{v}\right)=\min_{f\in\Pi(u,\mathrm{v})}<T,\text{ Cost }>_{F}$
-where $\mathcal{WD}(,)$ denotes the wasserstein distance, $\Pi\left(u,v\right)=\{T\in\mathbb{R}_{+}^{2C\times 2C}\mid T1_{2C}=u,T^{T}1_{2C}=v\}$, $1_{2C}$ represents a $2C$-dimensional all-one vector, $<,>_{F}$ refers to the Frobenius dot product, and $\text{Cost}\geq 0\in\mathbb{R}^{2C\times 2C}$ is a cost matrix calculated by the cosine distance between $x^{Phy}$ and $x^{v}$.
-The matrix $T\in\mathbb{R}^{2C\times 2C}$ is denoted as the optimal transport matching flow.
-Then, we apply the Pytorch framework-based batch-wise implementation of the Sinkhorn algorithm proposed in (Chen et al., 2020) to solve $\mathcal{WD}$.
-In this way, the obtained optimal transport matching flow $T$ can be used as a “cross-modal correlation” to enhance the multi-modal fusion.
-As such, the final multi-modal representation is adopted as follows:
+[…]
 (6) $x^{m}=(\text{Transformer}_{l}(x^{Phy}\otimes T)||\text{Transformer}_{l}(x^{v}))$
@@
 3.3. Label Correlation-driven Cross-Attention
-In the emotion distribution learning scenario, correlation across different labels plays a vital role in the prediction of emotional distributions.
-Therefore, we design a label correlation-driven cross-attention mechanism (LCDCA), shown in Figure 3, to effectively enhance the prediction performance of the model.
-Specifically, we introduce a learnable label embedding $x^{L}\in\mathbb{R}^{l\times d}$, where $l$ is the number of emotion classes.
-To learn the label correlation, we first generate learnable label correlation $M^{L}\in\mathbb{R}^{l\times l}$ and ground-truth label correlation $M^{gt}\in\mathbb{R}^{l\times l}$ through the cosine similarity, whose $(i,j)$-th element is given as:
+[…]
 (7) $M_{i,j}^{L}=\frac{x_{i}^{L}\cdot x_{j}^{L}}{||x_{i}^{L}||\cdot||x_{j}^{L}||},M_{i,j}^{gt}=\frac{L_{i}\cdot L_{j}}{||L_{i}||\cdot||L_{j}||}$
@@
 The boldfaced scores denote the best performances and the underlined scores are the second runners.
-The results of the proposed HeLo are presented in a gray background.
+The results of the proposed HeLo are presented in a blue background.
 Dataset Measure TKDE16 (Geng, 2016) TKDE16 (Geng, 2016) TPAMI13 (Geng et al., 2013) TNSRE22 (Song et al., 2022) ACMMM23 (Jiang et al., 2023) AAAI24 (Peng et al., 2024) TKDE23 (Jia et al., 2023) IJCAI24 (Kou et al., 2024) ICCV23 (Wen et al., 2023) TAFFC23 (Liu et al., 2023) Ours
@@
 The embedding size in the attention mechanism and feed-forward network is set to 128, and 64, respectively, and we set the depth of transformer blocks to 1.
-Our code will be released soon.
+Our code is released at: https://github.com/kaio-99/HeLo.
 4.4. Comparison with State-of-the-art Methods
@@
 Settings Measure w/o EEG w/o GSR w/o PPG w/o Video Ours
-Subject-Dependent Chebyshev (↓) 0.0529 0.0574 0.0487 0.0552 0.0446
+Subject-Dependent
+Chebyshev (↓) 0.0529 0.0574 0.0487 0.0552 0.0446
 Clark (↓) 0.3372 0.3550 0.3671 0.3380 0.3256
@@
 Intersection (↑) 0.9102 0.9007 0.9086 0.9033 0.9128
-Subject-Independent Chebyshev (↓) 0.0977 0.0953 0.0964 0.0926 0.0882
+Subject-Independent
+Chebyshev (↓) 0.0977 0.0953 0.0964 0.0926 0.0882
 Clark (↓) 0.6906 0.6742 0.6470 0.6635 0.6289
@@
 Settings Measure w/o ECG w/o EMG w/o EDA w/o ACC Ours
-Subject-Dependent Chebyshev (↓) 0.0088 0.0074 0.0079 0.0082 0.0073
+Subject-Dependent
+Chebyshev (↓) 0.0088 0.0074 0.0079 0.0082 0.0073
 Clark (↓) 0.0744 0.0722 0.0762 0.0695 0.0653
@@
 Intersection (↑) 0.9873 0.9899 0.9892 0.9885 0.9905
-Subject-Independent Chebyshev (↓) 0.0502 0.0419 0.0429 0.0474 0.0403
+Subject-Independent
+Chebyshev (↓) 0.0502 0.0419 0.0429 0.0474 0.0403
 Clark (↓) 0.3877 0.3597 0.3721 0.3842 0.3455
@@
 Intersection (↑) 0.8870 0.9095 0.9017 0.8991 0.9154
-Table 2 illustrates the subject-independent evaluation of the comparison methods and our proposed model on the DMER and WESAD datasets.
+Table 2 illustrates the subject-independent evaluation of the comparison methods and our proposed model on the DMER datasets.
 Except for the KL metric on the WESAD dataset, our method achieves the best results across both datasets and shows the highest average rank.
@@
 (d) WESAD Subject-Independent
-Figure 5. Ablation of network components in the proposed method. The experiment is conducted on the DMER and WESAD datasets, using both subject-dependent and subject-independent evaluation.
+Figure 5. Ablation of network components in HeLo.
 4.5. Ablation Studies
@@
 Ablation on the modalities.
-We also assessed the impact of each modality by separately removing each modality from our full model, the results are shown in Table 3 and Table 4.
+We also assessed the impact of each modality by separately removing each modality from our full model, the results are shown in Table 3 and 4.
 These experiments were conducted in both subject-dependent and subject-independent settings across both datasets.
 The results indicate that each modality contributes significantly to the overall performance of multi-modal emotion distribution learning, effectively integrating multiple modalities can enhance the emotion distribution performance.
-More ablation results on modalities can be referred to the supplementary materials.
-Figure 6. Effect of the number of attention heads. It shows the performance over six metrics on the DMER dataset.
-Figure 7. The effect of the depth of transformer block in the model on six evaluation metrics on the DMER dataset.
+Figure 6. Effect of attention heads on the DMER
+Figure 7. Effect of transformer block depth on the DMER.
 Ablation on the number of attention heads.
 The number of attention heads is usually an essential parameter for models based on attention mechanisms.
-Thus, we also compared the effect of different choices of the number of heads on the six evaluation metrics of the model on the DMER dataset in both subject-dependent and subject-independent settings, shown in Figure 6.
-It can be seen that the model’s performance first increases across subject-dependent and subject-independent settings as the number of attention heads increases.
-Then, the model achieves the best performance across all six evaluation metrics when the number of attention heads is equal to 4, after which the performance declines as the number of heads continues to increase.
-However, the number of attention heads has a relatively limited impact on the performance of the model.
+The result in Figure 6 indicates that, the model’s performance in both subject-dependent and subject-independent settings initially improves as the number of heads increases, and subsequently declines.
+The model reaches its optimal performance across all six evaluation metrics when the number of attention heads was 4.
+However, the overall impact of the number of attention heads on the model’s performance is relatively limited.
 Ablation on the depth of transformer block.
 The depth of end-to-end models is considered a crucial factor in determining their fitting ability.
-Therefore, we evaluate the model while varying the depth of the block on the DMER dataset to explore the impact of the depth of the transformer block.
-Figure 7 shows the optimal classification performance was achieved when the depth of the block was 1, and as the depth increased, the classification accuracy tended to slightly decrease and leveled off as the depth continued to increase, as the greater the depth of the transformer block, the more likely it is to overfit the model.
+Figure 7 illustrates that, the model achieved optimal classification performance when the transformer block depth was 1.
+As the depth increased, the classification accuracy showed a slight tendency to decrease and then leveled off.
+This is because a greater transformer block depth might cause overfitting.
 Figure 8. Visualization of the feature embedding.
@@
 Visualization of the cross-modal representations.
-We visualize the multi-modal feature space in the testing set by utilizing t-SNE (Van der Maaten and Hinton, 2008), and the results are illustrated in Figure 8.
+We visualize the multi-modal feature space in the testing set by utilizing t-SNE (Van der Maaten and Hinton, 2008).
 In Figure 8 (a), before the OTHM module, there is a noticeable separation between physiological (blue) and behavioral (red) representations.
@@
 All of these conclusions are consistent with human intuition, so as to demonstrate the better emotion distribution learning performance of our proposed model.
-Table 5. Comparison between our method and other models used in our experiments in terms of network efficiency.
-Metric TNSRE22 (Song et al., 2022) ACMMM23 (Jiang et al., 2023) AAAI24 (Peng et al., 2024) TAFFC23 (Liu et al., 2023) Ours
-Conformer MAET CARAT EmotionDict
-FLOPs 2.422M 91.955M 330.671M 156.354M 11.656M
-Params 1.023M 15.672M 6.780M 0.483M 4.040M
-4.7. Model Efficiency
-We compared the efficiency of our model with 10 other methods.
-On one hand, deep learning-based methods generally achieve better emotion distribution learning performance than traditional label distribution learning methods.
-On the other hand, traditional label distribution learning typically have fewer parameters.
-Therefore, we conducted a comparison of FLOPs (Floating Point Operations per Second) and the number of parameters that need to be learned in the model with the other four deep learning-based methods, i.e., Conformer (Song et al., 2022), MAET (Jiang et al., 2023), CARAT (Peng et al., 2024), and EmotionDict (Liu et al., 2023)).
-The results are summarized in Table 5.
 5. Conclusion
@@
 Experimental results on two multi-modal emotion datasets demonstrate the efficacy of our proposed method.
-In the future, we plan to utilize LLMs to extract the label embeddings, which provide strong semantic priors to our model instead of random initialization.
 Supplementary Material
@@
 We conducted thorough ablation experiments on different combinations of different emotion modalities.
-The results are illustrated in Table 9, 10, 11, and 12.
+The results are illustrated in Table 8, 9, 10, and 11.
 The experimental results demonstrate the necessity and effectiveness of our proposed multi-modal emotion distribution learning framework.
+Table 12. Comparison between our method and other models used in our experiments in terms of network efficiency.
+Metric TNSRE22 (Song et al., 2022) ACMMM23 (Jiang et al., 2023) AAAI24 (Peng et al., 2024) TAFFC23 (Liu et al., 2023) Ours
+Conformer MAET CARAT EmotionDict
+FLOPs 2.422M 91.955M 330.671M 156.354M 11.656M
+Params 1.023M 15.672M 6.780M 0.483M 4.040M
+Appendix E Model Efficiency
+We compared the efficiency of our model with 10 other methods.
+On one hand, deep learning-based methods generally achieve better emotion distribution learning performance than traditional label distribution learning methods.
+On the other hand, traditional label distribution learning typically have fewer parameters.
+Therefore, we conducted a comparison of FLOPs (Floating Point Operations per Second) and the number of parameters that need to be learned in the model with the other four deep learning-based methods, i.e., Conformer (Song et al., 2022), MAET (Jiang et al., 2023), CARAT (Peng et al., 2024), and EmotionDict (Liu et al., 2023)).
+The results are summarized in Table 12.